### v1

# BPE: Behavioral Profiling Ensemble

Abstract
Ensemble learning is widely recognized as a pivotal strategy for pushing the boundaries of predictive performance. Traditional static ensemble methods, such as Stacking, typically assign weights by treating each base learner as a holistic entity, thereby overlooking the fact that individual models exhibit varying degrees of competence across different regions of the instance space. To address this limitation, Dynamic Ensemble Selection (DES) was introduced. However, both static and dynamic approaches predominantly rely on the divergence among different models as the basis for integration. This inter-model perspective neglects the intrinsic characteristics of the models themselves and necessitates a heavy reliance on validation sets for competence estimation. In this paper, we propose the Behavioral Profiling Ensemble (BPE) framework, which introduces a novel paradigm shift. Unlike traditional methods, BPE constructs a “behavioral profile” intrinsic to each model and derives integration weights based on the deviation between the model’s response to a specific test instance and its established behavioral profile. Extensive experiments on both synthetic and real-world datasets demonstrate that the algorithm derived from the BPE framework achieves significant improvements over state-of-the-art ensemble baselines. These gains are evident not only in predictive accuracy but also in computational efficiency and storage resource utilization across various scenarios.
keywords:
Ensemble Learning, Behavioral Profiling, Dynamic Ensemble Selection
journal: Nuclear Physics B
[…]
[inst2]organization=Yunnan University, city=Kunming, state=Yunnan, country=China
1 Introduction
In the field of machine learning, ensemble learning, which enhances predictive performance by combining multiple base learners, has become the dominant approach for addressing complex classification problems. Whether in data science competitions or high-stakes industrial applications such as medical diagnosis and financial risk control, ensemble models typically exhibit superior generalization capability and robustness compared to individual models by exploiting diversity among learners [1, 2]. Current mainstream ensemble strategies are generally categorized into two paradigms: global static ensembles, such as simple averaging, median averaging, or rank averaging [3, 4]; and meta-learning-based methods, such as Stacking, which learn the combination weights of base models by training a meta-classifier [5].
Despite their substantial success, these methods predominantly rely on the assumption that a specific model possesses consistent superiority across the entire data distribution, thereby treating models as globally uniform. However, in practical scenarios, different models often manifest distinct strengths in various sample subspaces. This global perspective tends to overlook the local competence variances of the models. To address this issue, Dynamic Ensemble Selection (DES) has emerged [6]. Conventional DES methods typically follow a neighborhood-based validation paradigm: during the inference phase, they retrieve similar samples from a validation set and dynamically assign weights based on the models’ historical performance on these local neighbors [7, 8]. Nevertheless, such reliance on an external validation set poses challenges in deployment and limits generalization in certain contexts.
To overcome these limitations, we propose a fundamental shift in perspective: transitioning from external validation matching to the exploitation of models’ intrinsic properties. This paradigm shift can be understood through an intuitive analogy: traditional DES algorithms resemble ”resume screening”, when a new task (test sample) arrives, the system reviews historical records to find individuals who performed well on similar past tasks (validation neighbors). The drawback is that for novel or outlier instances, the historical resume loses its reference value. In contrast, our proposed method functions as ”behavioral profiling” rather than focusing on whether the individual has encountered similar tasks before, we observe their inherent behavioral traits through instantaneous stress testing to designate the most reliable candidate.
Driven by this motivation, this paper proposes the Behavioral Profiling Ensemble (BPE), a validation-free dynamic ensemble framework based on model ”personality” profiling. Specifically, we discard the traditional neighborhood search and instead introduce stochastic perturbations in the feature space to construct an instantaneous ”testing field” for each model. By observing the variations in the output distributions of different base models under perturbation (e.g., entropy increase), we can reveal the models’ inherent behavioral characteristics [9, 10]. This profiling effectively differentiates between models that are robustly confident and those that exhibit high uncertainty on the current sample, thereby serving as the basis for dynamic weighting. Extensive experiments on both simulation and real-world datasets demonstrate that BPE consistently outperforms traditional methods across numerous scenarios.
The remainder of this paper is organized as follows: Section 2 reviews related work in ensemble learning. Section 3 introduces the BPE framework and an entropy-based implementation algorithm. Section 4 presents comprehensive simulation experiments to verify the effectiveness of the algorithm. Section 5 evaluates the performance using a wide range of real-world datasets. Finally, Section 6 discusses future directions for the BPE framework.
2 Related Work
2.1 Ensemble Learning Theory and Uncertainty Measures
The core philosophy of ensemble learning lies in compensating for the inductive bias of a single learner by combining multiple hypotheses. The “No Free Lunch” theorem proposed by Wolpert and Macready [11] theoretically elucidates the necessity of multiple classifier systems, asserting that no single algorithm can perform optimally across all possible data distributions. To quantify the effectiveness of ensembles, Krogh and Vedelsby proposed the renowned “Error-Ambiguity Decomposition” theory [1], proving that the generalization error of an ensemble is strictly equal to the average error of the base models minus an ambiguity term representing model diversity. This theory establishes the dual objectives of ensemble learning: maximizing diversity among models while pursuing individual accuracy. From a statistical perspective, Dietterich further noted that ensemble methods effectively smooth decision boundaries by averaging multiple hypotheses, thereby significantly mitigating the dilemma of the bias-variance tradeoff for individual models [2].
To accurately allocate weights in dynamic ensembles, quantifying a single model’s “Predictive Confidence” on specific samples is crucial. Information entropy, as proposed by Shannon [12], serves as the cornerstone for measuring the uncertainty of probability distributions and is widely utilized to quantify classifier prediction confidence. Grandvalet and Bengio further demonstrated in semi-supervised learning that minimizing the entropy of the predictive distribution can effectively enhance model confidence on unlabeled data, establishing the theoretical connection between low entropy and high confidence [13]. In modern machine learning research, perturbation-based uncertainty estimation has become mainstream. Deep Ensembles, proposed by Lakshminarayanan et al. [10], demonstrated that epistemic uncertainty can be accurately estimated by ensembling multiple models. Similarly, Gal and Ghahramani demonstrated that enabling Dropout during the inference phase is equivalent to Bayesian approximation; by calculating predictive entropy through multiple stochastic forward passes, one can effectively capture the model’s intrinsic uncertainty [9]. Furthermore, Schapire’s margin theory indicates that maximizing the classification margin—manifested as significant differences in predictive probabilities—is key to enhancing generalization capability [14]. These studies provide a solid theoretical foundation for the proposed BPE framework, which constructs a behavioral profile of model entropy via training set perturbations.
2.2 Static Ensembles
Early ensemble strategies predominantly adopted a static perspective, where the performance weights of base models remain fixed across the entire feature space. Depending on the generation mechanism of the base learners, these strategies are primarily categorized into homogeneous and heterogeneous ensembles.
In homogeneous ensembles, the Bagging algorithm proposed by Breiman [15] constructs diversity via bootstrap sampling; its prominent representative, Random Forest [16], primarily enhances performance by mitigating variance. Conversely, Boosting paradigms aim to reduce bias by iteratively correcting residuals. Notable examples include the classic AdaBoost [17] and Gradient Boosting Machines (GBM) [18], along with highly optimized implementations such as XGBoost [19], LightGBM [20], and CatBoost [21].
In the realm of heterogeneous ensembles comprising diverse algorithms (e.g., SVC [22], KNN [23], and Logistic Regression [24]), the prevailing strategy involves static fusion rules. Kittler et al. systematized fundamental combination rules, including majority voting and simple averaging [3]. To address scale discrepancies in probability outputs, rank-based methods, such as the Borda Count applied by Ho et al. [4], mitigate these differences through rank aggregation. Furthermore, the Stacking framework [5] employs a meta-learner to learn the globally optimal combination of base models. Despite their widespread success, these static methods inherently overlook the locality of data distributions; a model with suboptimal global performance may typically exhibit superior competence within specific local subspaces.
2.3 Dynamic Ensembles
To address the rigidity of static ensembles, Dynamic Ensemble Selection (DES) has emerged [6]. Its core philosophy involves dynamically selecting the most competent subset of classifiers for each individual test sample. Existing DES methods primarily evaluate models based on a “Region of Competence” (RoC). Approaches predicated on local accuracy, such as LCA proposed by Woods et al. [7], assess models by computing classification accuracy within a local neighborhood. The KNORA series [8] adopts a more direct strategy, determining model participation based on the count of correctly classified samples in the vicinity. To circumvent the limitations of hard thresholds, RRC [25] utilizes Gaussian potential functions to assign weights based on the distance between test samples and neighbors, achieving a probabilistic competence assessment. MCB [26] introduces the concept of a “Behavioral Knowledge Space” to measure reliability by comparing the consistency of classifier output profiles in local regions. Meta-learning-based methods, such as META-DES [27], transform dynamic selection into a binary classification problem, utilizing meta-features to train a meta-classifier for competence estimation. Furthermore, FIRE-DES++ [28] incorporates online pruning strategies to enhance computational efficiency.
Despite their theoretical sophistication, these DES methods encounter significant challenges regarding their dependency on external reference sets for practical deployment. Fundamentally, existing DES methods adhere to a retrieval-based paradigm: the system is required to retain the entire reference set and execute computationally expensive nearest-neighbor searches during inference. As data scales expand, time and space complexity increase linearly or even super-linearly. Moreover, in high-dimensional feature spaces, as noted by Beyer et al. [29], the discriminative power of Euclidean distance diminishes sharply (the “curse of dimensionality”), resulting in retrieved “neighbors” that are often spurious and misleading for competence assessment. In summary, while DES offers a perspective of dynamic adaptation, its structural reliance on external reference sets and neighborhood retrieval constrains its scalability to larger and higher-dimensional datasets. This limitation motivates our exploration of a novel ensemble paradigm that achieves efficient and robust dynamic integration by constructing intrinsic behavioral profiles of the models.
3 BPE: Behavioral Profiling Ensemble
To address the limitations of conventional DES methods regarding storage overhead, retrieval latency, and failure in high-dimensional spaces, this section proposes BPE, a validation-free dynamic ensemble framework. The BPE framework abandons the traditional evaluation paradigm that relies on external validation sets for neighborhood retrieval. Instead, it adopts the concept of constructing model behavioral profiles via training set perturbations to reveal the intrinsic behavioral characteristics of base classifiers. Specifically, Section 3.1 elucidates the design philosophy of BPE, focusing on the paradigm shift from “resume screening” to “behavioral profiling.” Section 3.2 details a specific implementation algorithm, BPE-Entropy, based on information entropy. Finally, Section 3.3 provides a systematic analysis of the algorithmic complexity, demonstrating its advantages in computational efficiency and deployment flexibility.
3.1 Design Philosophy of BPE
The core design philosophy of the BPE framework lies in shifting the evaluation perspective of dynamic ensembles from external validation to internal consistency. Existing DES methods essentially follow a neighborhood-based validation paradigm, which assumes that the feature space surrounding a test sample shares similar decision boundaries with certain historical samples in the reference set. Under this assumption, the system infers a model’s performance on the current sample by retrieving neighbors from the reference set and calculating the model’s accuracy on these proxies.
However, this paradigm faces both theoretical and engineering challenges in practical applications. Theoretically, “distance” in high-dimensional feature spaces often loses its discriminative power [30], causing neighbors retrieved via Euclidean distance to be spurious (semantically irrelevant), thus introducing noisy evaluations. From an engineering perspective, reliance on a reference set necessitates maintaining an extensive historical database and performing expensive search computations during inference, which not only increases storage and latency overheads but also hinders the system’s ability to handle streaming or privacy-sensitive data.
To this end, we propose a novel paradigm centered on model behavioral profiling. We posit that the decision reliability of a model for a given instance should not be determined solely by its historical performance metrics, but rather by its intrinsic state when confronted with the current input. To capture these inherent behavioral attributes, we introduce training set perturbation as the core mechanism for establishing behavioral profiles. This approach is analogous to a “stress test” in psychology: rather than relying on a subject’s past records, we observe their authentic reactions under controlled environmental pressure.
Within the BPE framework, we construct a testing environment for profiling by injecting noise into the training set. Through this process, the divergence patterns in the model’s output distribution implicitly reveal its latent behavioral characteristics. Consequently, this allows us to discern when a model exhibits high uncertainty versus high predictive confidence. We allocate greater weight to a model when it demonstrates robust confidence. By calibrating a model’s response to different samples relative to its established behavioral profile, we derive a robust basis for dynamic weight assignment.
This approach offers several distinct advantages over existing methods:
  • 1.
    Minimal Storage Overhead: The maintenance cost is significantly lower than that of traditional DES methods, as BPE requires storing only a compact behavioral profile (typically a vector) for each model, rather than a raw reference dataset.
  • 2.
    Validation-Free & Data Efficiency: Since the construction of behavioral profiles does not mandate a separate validation set with ground-truth labels, the framework is highly effective in data-scarce regimes. It avoids the structural dependency on hold-out sets and eliminates the substantial computational costs associated with Out-Of-Fold (OOF) prediction techniques typically required to utilize full datasets.
  • 3.
    Inference Efficiency: Unlike retrieval-based DES methods, BPE necessitates only the calculation of lightweight statistical metrics during the inference phase. By eliminating complex nearest-neighbor search processes, BPE achieves a substantial advantage in inference speed, particularly in scenarios involving large-scale datasets.
Figure 1: Average Friedman ranks of BPE and twelve baseline models in homogeneous ensemble case.
3.2 BPE-Entropy
This section details the implementation of the BPE-Entropy algorithm. The algorithm utilizes the negative information entropy of the predictive distribution as the core metric for measuring intrinsic model uncertainty. By constructing behavioral profiles based on this metric, we facilitate dynamic weighting and achieve self-adaptive fusion of base classifiers.
Let the training set be denoted as $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, and let $\mathcal{M}=\{h_{1},\dots,h_{K}\}$ be a set of $K$ pre-trained heterogeneous base classifiers. For any input $\mathbf{x}$, the output of the $k$-th model regarding the $C$ classes is represented as a probability distribution vector $\mathbf{p}_{k}(\mathbf{x})=[p_{k,1},\dots,p_{k,C}]^{\top}$. To quantify the confidence of a model on a specific sample, we define the confidence score function $S(\cdot)$ as the negative information entropy of the predictive distribution:
$S(\mathbf{p}_{k}(\mathbf{x}))=\sum_{c=1}^{C}p_{k,c}(\mathbf{x})\log p_{k,c}(\mathbf{x})$ (1)
The use of negative entropy ensures a positive correlation between this metric and model confidence; that is, a larger $S$ value indicates a sharper predictive distribution and higher model certainty.
The first stage of the algorithm involves offline behavioral profiling, aimed at establishing a performance baseline for each model under simulated real-world conditions. Since profiling does not strictly rely on ground-truth labels, we move beyond traditional validation set retrieval and instead apply Gaussian perturbations to the training set feature space to simulate uncertainty in the testing environment.
This strategic choice is underpinned by critical advantages regarding efficiency and stability. Primarily, utilizing the training set allows for full-data utilization without incurring the computational overhead associated with Out-of-Fold (OOF) predictions, ensuring that base models are trained on the complete dataset. Furthermore, the training set provides a substantially larger sample size than a typical validation set ($N_{train}\gg N_{val}$), which generally yields a more reliable estimation of average model behavior and reduces the risk of skewing the profile due to small-sample randomness. Crucially, relying on a validation set carries the risk of extreme scenarios—where hold-out data is inadvertently too difficult or too trivial—which would distort the estimated baseline; in contrast, applying fixed Gaussian perturbation creates a standardized stress-test environment that eliminates such bias caused by the arbitrary difficulty of unseen samples.
To implement this simulation while ensuring mathematical consistency, we explicitly handle nominal features prior to perturbation. Specifically, all categorical attributes are transformed into numerical vectors via One-Hot Encoding, ensuring the input $\mathbf{x}_{i}$ resides in a continuous metric space. The noise injection is applied to this encoded space. The selection of a Gaussian distribution for the noise vector $\epsilon$ is theoretically grounded in the Principle of Maximum Entropy proposed by Jaynes [31]. According to this principle, the Gaussian distribution has the maximum entropy among all distributions with a specified variance. Choosing this distribution ensures that we introduce the least amount of structural bias into the simulation, as it avoids implicit assumptions (such as the hard boundaries inherent in a uniform distribution) that are unjustified without prior knowledge. Consequently, for each sample $\mathbf{x}_{i}$, in the training set, we generate a perturbed sample $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{i}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\delta^{2}\mathbf{I})$ represents the unbiased noise vector.
Subsequently, we characterize the distribution of confidence scores for each base model across the entire perturbed dataset by calculating the mean $\mu_{k}$ and standard deviation $\sigma_{k}$:
$\mu_{k}=\frac{1}{N}\sum_{i=1}^{N}S(\mathbf{p}_{k}(\tilde{\mathbf{x}}_{i})),\quad\sigma_{k}=\sqrt{\frac{1}{N-1}\sum_{i=1}^{N}(S(\mathbf{p}_{k}(\tilde{\mathbf{x}}_{i}))-\mu_{k})^{2}}$ (2)
We define the tuple $\mathcal{P}_{k}=(\mu_{k},\sigma_{k})$ as the behavioral profile of model $h_{k}$, where $\mu_{k}$ represents the average confidence level when facing uncertainty, and $\sigma_{k}$ characterizes the sensitivity of its confidence fluctuations. This process is executed only once post-training, with the parameters $\mathcal{P}_{k}$ stored as lightweight metadata.
The second stage of the algorithm is online dynamic weighting. During the inference phase, for a new test sample $\mathbf{x}_{\text{test}}$, the system bypasses neighborhood searches and directly computes the instantaneous confidence for each model:
$S_{\text{test},k}=S(\mathbf{p}_{k}(\mathbf{x}_{\text{test}}))$ (3)
To eliminate scale disparities in probabilities caused by differing algorithmic principles among base models, we introduce the concept of relative confidence. Utilizing the pre-stored profiling parameters, we normalize the instantaneous confidence via a Z-score transformation:
$z_{k}(\mathbf{x}_{\text{test}})=\frac{S_{\text{test},k}-\mu_{k}}{\sigma_{k}+\xi}$ (4)
where $\xi$ is a constant for numerical stability. The physical interpretation of $z_{k}$ is intuitive: it measures the number of standard deviations the current model confidence deviates from its historical behavioral profile. If $z_{k}>0$, the model exhibits higher-than-usual confidence and should be rewarded; otherwise, the model is considered confused and should be suppressed. Finally, we map these normalized scores to non-negative weights. To amplify the contribution of high-confidence models, we employ an exponential mapping function followed by normalization:
$w_{k}(\mathbf{x}_{\text{test}})=\frac{\exp(\lambda\cdot z_{k}(\mathbf{x}_{\text{test}}))}{\sum_{j=1}^{K}\exp(\lambda\cdot z_{j}(\mathbf{x}_{\text{test}}))}$ (5)
where $\lambda$ is a sensitivity hyperparameter. The final ensemble output $H(\mathbf{x}_{\text{test}})$ is derived from the weighted combination of the base models. Owing to the versatility of the BPE framework, these weights can be applied directly to probability vectors or to ranking results, flexibly adapting to different task requirements:
$H(\mathbf{x}_{\text{test}})=\sum_{k=1}^{K}w_{k}(\mathbf{x}_{\text{test}})\cdot\Phi(\mathbf{x}_{\text{test}})_{k}$ (6)
where $\Phi(\cdot)_{k}$ represents the output format of the $k$-th model (e.g., probability vector or rank vector).
Algorithm 1 The BPE-Entropy Algorithm
1:Training dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$; Pool of base classifiers $\mathcal{M}=\{h_{1},\dots,h_{K}\}$; Perturbation scale $\delta$; Sensitivity factor $\lambda$.
2:Ensemble prediction $H(\mathbf{x}_{\text{test}})$.
3:// Phase 1: Offline Profiling
4:Generate perturbed dataset $\tilde{\mathcal{D}}$ by adding Gaussian noise $\epsilon\sim\mathcal{N}(0,\delta^{2}\mathbf{I})$ to $\mathcal{D}$.
5:for $k=1$ to $K$ do
6:  Compute confidence scores (negative entropy) for all samples in $\tilde{\mathcal{D}}$ using $h_{k}$.
7:  Calculate mean $\mu_{k}$ and standard deviation $\sigma_{k}$ of the scores.
8:  Store behavioral profile $\mathcal{P}_{k}\leftarrow(\mu_{k},\sigma_{k})$.
9:end for
10:// Phase 2: Online Dynamic Weighting
11:Input: New test sample $\mathbf{x}_{\text{test}}$.
12:for $k=1$ to $K$ do
13:  Compute instantaneous confidence: $S_{k}\leftarrow\text{NegEntropy}(h_{k}(\mathbf{x}_{\text{test}}))$.
14:  Compute Z-score standardization:
15:   $z_{k}\leftarrow(S_{k}-\mu_{k})/(\sigma_{k}+\xi)$
16:  Compute unnormalized weight (Exponential Mapping):
17:   $w^{\prime}_{k}\leftarrow\exp(\lambda\cdot z_{k})$
18:end for
19:Normalize weights: $w_{k}\leftarrow w^{\prime}_{k}/\sum_{j=1}^{K}w^{\prime}_{j}$.
20:Return Ensemble output $H(\mathbf{x}_{\text{test}})=\sum_{k=1}^{K}w_{k}\cdot\Phi(\mathbf{x}_{\text{test}})_{k}$.
3.3 Complexity Analysis
To comprehensively evaluate the deployment potential of the BPE framework, this section analyzes its complexity in terms of space complexity and time complexity, contrasting it with conventional neighborhood-based DES approaches such as KNORA [8], LCA [7], and RRC [25].
Traditional DES methods necessitate maintaining the entire reference set (validation or training set) during the inference phase for retrieval purposes. Assuming a reference set size of $N$, a feature dimensionality of $D$, and $K$ base models, the space complexity of traditional DES is primarily dictated by the scale of the reference set, approximating $\mathcal{O}(N\cdot D+N\cdot K)$. In large-scale data scenarios, memory consumption grows linearly with sample accumulation, imposing significant constraints on storage-limited devices. In contrast, BPE compresses the distributional characteristics into concise “behavioral profiles” during the offline phase. For each base model, BPE solely requires storing two scalar statistics ($\mu_{k}$ and $\sigma_{k}$), resulting in an auxiliary storage overhead of merely $\mathcal{O}(K)$. Given that $K\ll N\cdot D$, BPE substantially lowers the barriers to deployment.
Regarding time complexity, it is essential to distinguish between offline profiling and online inference. In the offline phase, both traditional DES and BPE require a comprehensive inference pass over the dataset to compute statistical features. However, during the online inference phase, traditional DES must execute a $k$-nearest neighbor search for each test sample. Even with optimized indexing structures like KD-Trees, search efficiency in high-dimensional spaces degrades towards $\mathcal{O}(N\cdot D)$, causing inference latency to escalate significantly with the volume of historical data.
Conversely, BPE completely eliminates the dependence on historical data during inference, requiring only the computation of the current sample’s entropy and Z-score. The complexity of computing entropy is $\mathcal{O}(C)$ (where $C$ is the number of classes), and the complexity of normalization and weighting is $\mathcal{O}(K)$. Consequently, the total online selection complexity of BPE remains constant at $\mathcal{O}(K\cdot C)$. This implies that the inference speed of BPE is entirely independent of the scale of the training or reference data, maintaining constant computational overhead regardless of the magnitude of historical data.
Table 1: Complexity comparison between traditional KNN-based DES and the proposed BPE framework. Note that $N$, $D$, $K$, and $C$ denote the dataset size, feature dimension, number of models, and number of classes, respectively.
Method Storage Cost (Space) Offline Preparation (Time) Online Inference (Time) Dependency on Dataset Size ($N$)
Traditional DES High: $\mathcal{O}(N\cdot D)$ High: $\mathcal{O}(N\cdot K)$ High: $\mathcal{O}(N\cdot D)$ Linear
BPE (Ours) Low: $\mathcal{O}(K)$ High: $\mathcal{O}(N\cdot K)$ Fast: $\mathcal{O}(1)$ None (Constant)
4 Simulation Experiments
4.1 Experimental Setup
4.1.1 Data Generation
To systematically evaluate the performance of BPE across varying dimensionalities and data characteristics, we designed three synthetic experimental scenarios. All experiments utilize Monte Carlo simulation methods, with each scenario executing 1,000 randomized sampling iterations within specified parameter ranges to ensure the statistical robustness of the results.
Scenario 1: Binary Gaussian Mixture Classification. This scenario evaluates the algorithm’s performance in fundamental classification tasks. Data are generated using a multi-dimensional Gaussian Mixture Model (GMM). For each class $c\in\{0,1\}$, samples $\mathbf{x}$ are drawn from a mixture of $K$ Gaussian components:
$\mathbf{x}\sim\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$ (7)
where $K$ represents the number of components, and $\boldsymbol{\mu}_{k}$ and $\boldsymbol{\Sigma}_{k}$ denote the mean vector and covariance matrix of the $k$-th component, respectively. To simulate observational noise inherent in real-world settings, we introduce label-flip noise into the training set. Let $y$ be the ground-truth label and $\tilde{y}$ be the observed label; the noise injection follows a Bernoulli distribution:
$P(\tilde{y}=y)=1-\rho,\quad P(\tilde{y}\neq y)=\rho$ (8)
In this experiment, the noise rate $\rho$ is sampled uniformly from the interval $[0.03,0.05]$, and the feature dimensionality $D$ is sampled uniformly from the integer set $\{10,\dots,30\}$.
Scenario 2: Multi-class Adaptability. This scenario examines the algorithm’s adaptability in complex multi-class environments. The data generation mechanism remains consistent with Scenario 1, with the exception that the number of classes $C$ is not fixed but follows a discrete uniform distribution $C\sim U\{3,10\}$. As $C$ increases, the overlapping regions in the feature space expand significantly, resulting in more intricate decision boundaries. This necessitates that the ensemble algorithm not only discriminates between positive and negative samples but also accurately quantifies the inter-class ambiguity via entropy.
Scenario 3: Non-linear Hastie Task. To verify the capability of the algorithm in handling non-linear manifolds, we employ the standard Hastie 10-2 dataset [30]. The decision boundary of this dataset is defined by a hypersphere and is linearly inseparable. For a 10-dimensional input vector $\mathbf{x}=(x_{1},\dots,x_{10})$, the label $y$ is determined by the following discriminant function:
$y=\begin{cases}1&\text{if }\sum_{j=1}^{10}x_{j}^{2}>\chi^{2}_{10}(0.5)\approx 9.34\\ 0&\text{otherwise}\end{cases}$ (9)
where $\chi^{2}_{10}(0.5)$ denotes the median of a chi-square distribution with 10 degrees of freedom. The concentric “onion-like” structure of this task presents a significant challenge to algorithms predicated on local linearity assumptions.
4.1.2 Base Model Pool
We assembled a diverse repository containing 13 heterogeneous base classifiers, spanning the mainstream algorithmic paradigms in machine learning:
  • 1.
    Bagging-based Models: Including Random Forest [16] and Extra Trees [32]. These methods mitigate the variance of individual decision trees through random feature subspaces and bootstrap aggregation.
  • 2.
    Boosting-based Models: Including XGBoost [19], LightGBM [20], CatBoost [21], and AdaBoost [17]. These models excel in modeling complex non-linear boundaries by iteratively optimizing residuals.
  • 3.
    Linear and Discriminant Models: Including Logistic Regression [24] and Linear Discriminant Analysis (LDA) [33], which provide robust linear approximations to prevent overfitting.
  • 4.
    Non-linear and Distance-based Models: Including MLPClassifier [34], SVC (with probability estimation) [22], and K-Neighbors (KNN) [23].
  • 5.
    Probabilistic Models: Gaussian Naive Bayes [35] is included to provide a probabilistic baseline.
To ensure experimental validity, we implemented a strict performance-based pruning mechanism. We retain only those base models $h_{k}$ whose error rate $E(h_{k})$ satisfies:
$E(h_{k})\leq E_{\text{best}}\times(1+\alpha)$ (10)
where $E_{\text{best}}$ denotes the error of the best-performing model on the validation set, and the tolerance threshold is set to $\alpha=0.15$. This mechanism effectively filters out suboptimal models (“weak links”) that are ill-suited for the current data distribution[36].
It should be noted that this setup is only for the purpose of the automated base model pool, because in actual deployment we do not include models with too large a gap in the integration. The validation set here is only used to simulate this situation.
To guarantee the fairness of the experiments, we adopted standard parameter settings for all base learners, deliberately refraining from specific hyperparameter tuning. This protocol ensures that the observed performance gains are attributable solely to the efficacy of the ensemble integration rather than individual model optimization. The specific hyperparameter configurations for all base learners are detailed in Table 2.
4.1.3 Baseline Methods
To rigorously evaluate the performance of BPE, we compare it against a comprehensive set of competitive ensemble strategies, categorized into static and dynamic approaches.
Static Ensemble Baselines.
  • 1.
    Simple Average: This method operates on the assumption that all base models contribute equally. It computes the arithmetic mean of the output probability vectors from all base classifiers as the final prediction. Despite its simplicity, consistent with Dietterich’s statistical perspective, simple averaging effectively smooths decision boundaries and reduces variance by aggregating multiple independent hypotheses, serving as a robust baseline that is often challenging to outperform significantly.
  • 2.
    Weighted Average: An empirical strategy derived from engineering practice. In industrial applications and data science competitions, the Out-of-Fold (OOF) accuracy generated during the cross-validation phase is typically utilized as static weights. This method relies on the intuitive assumption that models demonstrating superior performance on the validation set should be granted higher authority during the testing phase.
  • 3.
    Rank Average: A robust aggregation technique widely adopted to address probability calibration discrepancies in heterogeneous ensembles. To mitigate the interference caused by scale inconsistencies across different algorithms, Rank Average discards absolute probability values in favor of fusing the relative rankings of predicted scores. It is widely regarded as a highly stable engineering baseline.
  • 4.
    Median Average: In scenarios where output probability vectors exhibit extreme outliers, the median offers superior robustness compared to the mean and serves as a commonly utilized ensemble technique.
  • 5.
    Stacking: A meta-learning-based ensemble framework proposed by Wolpert [5]. In this study, logistic regression is employed as the meta-learner. The training process utilizes the OOF predictions of base models on the validation set as meta-features and ground-truth labels as targets to learn a set of linear combination weights that minimize global loss. While Stacking can automatically capture the complementarity between models, its learned weights remain fixed during inference and cannot adapt dynamically to individual test samples.
Table 2: Hyperparameter configurations for the base model pool.
Category Algorithm Specific Configurations / Key Parameters
Bagging Random Forest n_estimators=100n\_estimators=100, $criterion=^{\prime}gini^{\prime}$
Extra Trees n_estimators=100n\_estimators=100, $bootstrap=False$
Boosting XGBoost $eval\_metric=^{\prime}logloss^{\prime}$, use_label_encoder=Falseuse\_label\_encoder=False
LightGBM verbose=1verbose=-1
CatBoost $logging\_level=^{\prime}Silent^{\prime}$, thread_count=1thread\_count=1
Gradient Boosting n_estimators=100n\_estimators=100, learning_rate=0.1learning\_rate=0.1
AdaBoost n_estimators=50n\_estimators=50, $algorithm=^{\prime}SAMME.R^{\prime}$
Linear Logistic Regression max_iter=1000max\_iter=1000, $solver=^{\prime}lbfgs^{\prime}$
LDA Default (SVD solver)
Others MLP Classifier max_iter=3000max\_iter=3000, early_stopping=Trueearly\_stopping=True
KNN n_neighbors=5n\_neighbors=5, $weights=^{\prime}uniform^{\prime}$
SVC $probability=True$, $kernel=^{\prime}rbf^{\prime}$
Gaussian NB Default
Dynamic Ensemble Baselines.
  • 1.
    LCA (Local Class Accuracy): A classic dynamic selection method proposed by Woods et al. [7]. It assigns weights based on local class accuracy. For each test sample, LCA retrieves the $k$-nearest neighbors from the reference set and calculates the classification accuracy of each base model regarding the predicted class within this local neighborhood. If a model demonstrates high historical local accuracy, it is deemed an “expert” for that specific sample.
  • 2.
    KNORA-Union (k-Nearest Oracles Union): Proposed by Ko et al. [8], this algorithm is predicated on the “Nearest Oracle” assumption. Unlike LCA, it employs a direct voting mechanism: for each neighbor of the test sample, every base model that correctly classifies that neighbor is granted a vote. This relatively lenient strategy tends to involve a broader subset of models in the decision-making process, thereby reducing predictive variance.
  • 3.
    KNORA-Eliminate (k-Nearest Oracles Eliminate): A stricter variant within the KNORA family. It mandates that a base model must correctly classify all $k$ neighbors surrounding the test sample to be eligible for voting. If no model satisfies this unanimous condition, the value of $k$ is iteratively decremented until a qualifying subset is identified.
  • 4.
    RRC (Randomized Reference Classifier): Proposed by Woloszynski [25] to mitigate the hard-thresholding limitations of traditional $k$-NN methods. RRC introduces a continuous probabilistic evaluation perspective by constructing a randomized reference classifier framework. It utilizes Gaussian kernels as potential functions to assign varying weights based on the Euclidean distance between reference and test samples, effectively smoothing the evaluation of local competence.
  • 5.
    MCB (Multiple Classifier Behavior): Proposed by Giacinto et al. [26], this method performs selection by analyzing the behavioral consistency of base models. Unlike methods focused solely on accuracy, MCB retrieves neighbors in the reference set that exhibit similar “output profiles” (vectors composed of all base model predictions) to the current test sample. The premise is that if historical output patterns match the current pattern, the labels of those neighbors should guide the final decision.
  • 6.
    DES-AS: A state-of-the-art synergy-based framework proposed by Zhang et al. [37]. It utilizes Algorithm Shapley, a game-theoretic variant of the Shapley Value, to measure the synergy effect among classifiers. By calculating the marginal contribution of each model to all possible ensemble permutations, it selects classifiers with positive synergy values for dynamic aggregation.
4.2 Basic Simulation Results
The performance of the proposed BPE framework and the various baseline methods across the three synthetic scenarios is summarized in Table 3. The results indicate that the BPE framework consistently achieves superior average accuracy (ACC) in Scenario 1 and Scenario 2 compared to the vast majority of baseline methods, with the sole exception of Stacking. It is worth noting that while Stacking achieves marginal gains in these linear-dominant scenarios, it entails significantly higher computational resource consumption due to meta-model training, whereas BPE remains validation-free and lightweight.
Statistical significance tests further validate the superiority of our approach. In comparison with the second-best performing algorithm (excluding Stacking) in each scenario, the p-values for BPE were $2.79\times 10^{-12}$ for Scenario 1 and $4.43\times 10^{-2}$ for Scenario 2, respectively.
Most notably, in Scenario 3, which involves a complex non-linear manifold (the Hastie hypersphere), the BPE algorithm demonstrates the most outstanding performance among all tested methods (including Stacking), achieving a statistical significance p-value of $1.15\times 10^{-10}$ against the runner-up. These results underscore the robustness and efficacy of the BPE framework, particularly its ability to capture intrinsic model behaviors in complex decision boundaries where traditional distance-based or linear meta-learning methods may falter.
Table 3: Classification accuracy comparison (mean of 1,000 runs) across three simulation scenarios. The best results are highlighted in bold-face.
Method Scenario 1 Scenario 2 Scenario 3
(Gaussian Binary) (Gaussian Multi-class) (Hastie Non-linear)
Static Baselines
Best Single 0.8932 0.6685 0.9106
Simple Avg 0.8996 0.6839 0.9111
Weighted Avg 0.8996 0.6846 0.9112
Rank Avg 0.8993 0.6877 0.9130
Stacking 0.9018 0.6916 0.9115
DES Baselines
LCA 0.8998 0.6836 0.9112
KNORA-U 0.8997 0.6822 0.9113
KNORA-E 0.8975 0.6826 0.9099
RRC 0.8998 0.6836 0.9112
MCB 0.8978 0.6822 0.9099
DES-AS 0.8993 0.6804 0.9112
Proposed Method
BPE-Entropy 0.9006 0.6841 0.9131
BPE-Rank 0.9001 0.6883 0.9136
4.3 Hyperparameter Sensitivity Analysis
As illustrated in Figure 2, the BPE framework exhibits a high degree of stability with respect to hyperparameter variations. With the minor exception of Scenario 2, where performance shows slight responsiveness to changes in the Noise Scale, the classification accuracy (ACC) remains remarkably stable across the majority of configurations.
To rigorously demonstrate the inherent robustness of the BPE algorithm, we deliberately refrained from performing any specific hyperparameter tuning for the experiments reported in this paper. All experimental results were derived using a fixed, unified configuration: a Sensitivity factor of $\lambda=1.0$ and a Noise Scale of $\delta=0.5$.
Figure 2: Hyperparameter sensitivity analysis. The heatmaps illustrate the impact of varying the Sensitivity factor ($\lambda$) and Noise Scale ($\delta$) on classification accuracy across the three simulated scenarios. The consistent performance across a wide range of parameters demonstrates the robustness of the BPE framework.
5 Real Data
5.1 Setting
We selected 40 real-world datasets from OpenML for experimentation[38]. These datasets originate from diverse high-quality sources, including the UCI Machine Learning Repository and NASA, covering a broad spectrum of domains ranging from software engineering (software defect prediction) and healthcare (disease diagnosis and physiological monitoring) to natural science (high-energy physics, biological classification, and climate simulation), business finance (customer churn prediction and credit approval), as well as cybersecurity and information technology (spam and phishing website detection). Detailed specifications of these datasets are provided in Table 6. The configuration of the base model pool remains consistent with that of the simulation experiments. Furthermore, we conducted homogeneous model ensemble experiments; although such scenarios typically manifest as encapsulated algorithms in practical deployment (e.g., Random Forest), our homogeneous ensemble comprises 40 individual Decision Tree models, where the selection protocol for integration remains identical to that of the heterogeneous ensemble.
Table 4: Summary description of the 40 datasets used in the experiments.
Dataset ID ID Dataset #Ex. #Atts. #Num. #Nom. #Cl.
D1 2 anneal 898 39 6 33 5
D2 7 audiology 226 70 0 70 24
D3 9 autos 205 26 15 11 6
D4 39 ecoli 336 8 7 1 8
D5 42 soybean 683 36 0 36 19
D6 43 haberman 306 4 2 2 2
D7 49 heart-c 303 14 6 8 2
D8 59 ionosphere 351 35 34 1 2
D9 62 zoo 101 17 1 16 7
D10 333 monks-problems-1 556 7 0 7 2
D11 337 SPECTF 349 45 44 1 2
D12 1017 arrhythmia 452 280 206 74 2
D13 1036 sylva_agnostic 14395 217 216 1 2
D14 1040 sylva_prior 14395 109 108 1 2
D15 1049 pc4 1458 38 37 1 2
D16 1050 pc3 1563 38 37 1 2
D17 1053 jm1 10885 22 21 1 2
D18 1063 kc2 522 22 21 1 2
D19 1067 kc1 2109 22 21 1 2
D20 1460 banana 5300 3 2 1 2
D21 1461 bank-marketing 45211 17 7 10 2
D22 1463 blogger 100 6 0 6 2
D23 1465 breast-tissue 106 10 9 1 6
D24 1466 cardiotocography 2126 36 35 1 10
D25 1467 Climate-model 540 21 20 1 2
D26 1480 ilpd 583 11 9 2 2
D27 1489 phoneme 5404 6 5 1 2
D28 1494 qsar-biodeg 1055 42 41 1 2
D29 1503 spoken-arabic-digit 263256 15 14 1 10
D30 1511 wholesale-customers 440 9 7 2 2
D31 1567 poker-hand 1025009 11 10 1 10
D32 4538 GesturePhase 9873 33 32 1 5
D33 6332 cylinder-bands 540 40 18 22 2
D34 40693 xd6 973 10 0 10 2
D35 40705 tokyo1 959 45 42 3 2
D36 40945 Titanic 1309 14 6 3 2
D37 40966 MiceProtein 1080 82 77 5 8
D38 40978 Internet-Ad 3279 1559 3 1556 2
D39 40981 Australian 690 15 6 9 2
D40 41144 madeline 3140 260 259 1 2
** For datasets with more than 10,000 samples, considering computational resource limitations, we uniformly downsample to 10,000 samples. Since we repeat the experiment 50 times with different random seeds, most samples from each dataset will actually be selected.
5.2 Heterogeneous
In the heterogeneous ensemble scenario, the proposed BPE framework demonstrates robust superiority over state-of-the-art baselines. As detailed in Table 5, BPE achieves the highest average classification accuracy of 83.06%, outperforming the second-best method, Stacking (82.93%), and significantly surpassing the best single classifier (82.83%). The stability of the proposed method is further corroborated by the Friedman ranking analysis presented in Fig. 3. BPE secures the lowest (best) average rank of 3.325, establishing a substantial lead over the runner-up, RRC, which has an average rank of 5.075.
To validate the statistical significance of these improvements, the Wilcoxon signed-rank test was conducted at a significance level of $\alpha=0.05$. The results, summarized in Table 6, indicate that the null hypothesis is rejected for 11 out of 12 comparisons, confirming that BPE performs statistically better than the majority of baseline models. The only exception is Stacking ($p=0.109$), where no significant statistical difference was observed. However, it is important to note that Stacking relies on training a complex meta-classifier, often requiring significantly higher computational resources. In contrast, BPE achieves comparable predictive performance through a more efficient behavioral profiling mechanism.
Table 5: Average classification accuracy obtained by BPE and twelve baseline models over 40 datasets in heterogeneous ensemble case. The best results are highlighted in bold-face.
Dataset Static DES Proposed
Best Simple Median Rank Stacking Weighted DES-AS KNE KNU LCA MCB RRC BPE
D1 98.59 98.62 98.63 63.77 98.59 98.62 98.62 98.66 98.61 98.64 98.65 98.65 98.66
D2 34.91 34.41 34.53 31.03 34.26 34.59 34.62 34.56 34.71 34.65 34.65 34.68 34.62
D3 79.19 79.10 79.16 73.74 78.65 79.16 78.81 79.16 78.94 79.00 78.87 78.90 79.52
D4 79.74 80.26 80.32 67.94 80.02 80.34 80.20 80.16 80.20 80.22 80.06 80.22 80.12
D5 93.20 93.20 93.22 87.27 93.29 93.19 93.25 93.29 93.24 93.23 93.30 93.22 93.45
D6 72.87 72.54 72.61 61.57 72.11 72.59 72.26 72.61 72.48 72.52 72.37 72.46 72.35
D7 81.87 82.24 82.29 81.60 82.00 82.33 82.33 82.13 82.31 82.35 82.11 82.31 82.15
D8 93.19 92.81 92.91 86.42 93.49 92.83 93.00 92.94 92.96 92.89 92.96 92.91 93.15
D9 93.55 93.68 93.61 89.42 91.48 93.74 93.68 93.42 93.61 93.61 93.48 93.61 93.16
D10 98.99 98.99 98.99 97.16 99.01 98.99 98.99 98.99 99.01 98.99 98.99 98.99 98.99
D11 87.92 88.10 87.79 77.01 88.72 88.11 87.64 88.06 87.83 88.02 87.45 88.02 88.06
D12 80.62 80.76 81.13 80.35 81.50 80.79 80.66 80.85 80.69 80.75 80.41 80.75 80.82
D13 99.35 99.42 99.42 65.11 99.41 99.42 99.41 99.40 99.42 99.42 99.41 99.42 99.43
D14 99.39 99.46 99.45 65.49 99.44 99.46 99.45 99.44 99.46 99.46 99.44 99.46 99.46
D15 90.65 91.18 91.20 64.67 91.14 91.19 91.24 91.08 91.18 91.23 91.09 91.23 91.16
D16 89.70 89.99 89.91 62.42 89.96 89.98 89.83 89.93 89.88 89.99 89.85 89.98 90.07
D17 81.26 81.50 81.43 63.10 81.77 81.50 81.68 81.63 81.72 81.66 81.62 81.72 81.90
D18 83.38 83.63 83.77 71.41 83.50 83.62 83.77 83.64 83.59 83.62 83.44 83.68 83.63
D19 85.54 86.47 86.30 64.01 86.41 86.47 86.43 86.50 86.48 86.45 86.28 86.54 86.54
D20 90.23 90.06 90.04 88.53 90.19 90.06 89.98 90.03 90.00 90.01 89.95 90.01 90.12
D21 90.25 90.36 90.37 64.48 90.38 90.36 90.32 90.30 90.34 90.34 90.26 90.34 90.37
D22 81.87 81.60 81.73 76.20 81.00 81.53 81.33 81.60 81.00 81.33 80.80 81.53 82.07
D23 40.25 42.13 41.81 43.25 39.44 42.50 42.00 42.06 42.38 42.50 42.31 42.63 42.31
D24 100.00 100.00 100.00 99.99 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
D25 91.01 91.01 91.02 59.83 91.14 91.01 90.95 90.99 91.02 91.00 90.93 91.00 91.04
D26 70.82 71.05 70.96 64.85 71.42 71.09 70.96 70.93 71.29 71.31 71.01 71.29 71.44
D27 90.64 90.45 90.46 78.90 90.82 90.46 90.43 90.47 90.48 90.45 90.44 90.45 90.58
D28 87.04 87.59 87.41 80.84 87.55 87.56 87.62 87.53 87.59 87.55 87.39 87.56 87.62
D29 11.35 11.32 10.99 10.94 10.41 11.32 11.08 11.32 11.02 11.09 11.09 11.10 11.42
D30 90.76 91.55 91.42 83.71 91.05 91.50 91.33 91.33 91.41 91.52 91.23 91.48 91.58
D31 42.19 41.05 41.77 27.36 40.13 41.21 41.87 41.05 41.89 41.68 41.66 41.69 40.90
D32 67.84 68.07 68.08 63.70 70.26 68.08 68.29 68.09 68.30 68.19 68.34 68.20 68.40
D33 79.99 79.63 79.74 78.31 81.33 79.78 79.99 79.74 80.05 79.95 79.35 79.95 80.31
D34 100.00 100.00 100.00 87.64 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00 100.00
D35 92.84 92.80 92.78 86.73 93.02 92.79 92.74 92.74 92.73 92.73 92.68 92.74 92.88
D36 80.63 81.05 81.11 77.31 81.10 81.05 80.96 80.93 81.03 81.05 80.81 81.04 81.19
D37 99.45 99.59 99.59 97.55 99.60 99.59 99.65 99.62 99.64 99.61 99.64 99.63 99.66
D38 97.72 97.76 97.73 72.59 97.85 97.76 97.76 97.78 97.78 97.77 97.78 97.78 97.82
D39 85.86 86.58 86.42 86.10 86.57 86.58 86.54 86.60 86.47 86.46 86.46 86.46 86.70
D40 81.59 82.21 82.17 82.25 82.37 82.21 82.23 81.99 82.22 82.23 81.89 82.23 82.27
Average 82.41 82.56 82.56 71.61 82.51 82.58 82.55 82.54 82.57 82.59 82.46 82.60 82.65
Table 6: Wilcoxon signed-rank test results for comparing BPE with twelve baseline models in heterogeneous ensemble case.
Ensemble type Comparison $^{+} RR^{-}$ Hypothesis $p$-value
Static BPE vs. Best Single 582.0 121.0 Rejected for BPE at 5% 0.0005**
Static BPE vs. Median Avg 526.0 177.0 Rejected for BPE at 5% 0.0085**
Static BPE vs. Rank Avg 813.0 7.0 Rejected for BPE at 5% 0.0000**
Static BPE vs. Simple Avg 527.0 139.0 Rejected for BPE at 5% 0.0023**
Static BPE vs. Stacking 490.0 251.0 Not Rejected 0.0831
Static BPE vs. Weighted Avg 517.0 186.0 Rejected for BPE at 5% 0.0125**
DES BPE vs. DES-AS 510.0 120.0 Rejected for BPE at 5% 0.0014**
DES BPE vs. KNE 543.0 87.0 Rejected for BPE at 5% 0.0002**
DES BPE vs. KNU 573.0 168.0 Rejected for BPE at 5% 0.0033**
DES BPE vs. LCA 521.0 182.0 Rejected for BPE at 5% 0.0106**
DES BPE vs. MCB 597.0 69.0 Rejected for BPE at 5% 0.0000**
DES BPE vs. RRC 500.0 203.0 Rejected for BPE at 5% 0.0251**
** Close to the $p$-value means that statistical differences are found with $\alpha=0.05$.
Figure 3: Average Friedman ranks of BPE and twelve baseline models in heterogeneous ensemble case.
Figure 4: Win-tie-loss distribution of the average classification accuracy of the proposed BPE vs. twelve baseline models over 40 datasets in heterogeneous ensemble scenario. Each full line illustrates the critical value $n_{c}=\{24.05,25.20,27.36\}$ considering confidence level of $\alpha=\{0.10,0.05,0.01\}$.
5.3 Homogeneous
In the homogeneous ensemble scenario, where diversity is generated solely through data resampling (Bagging), BPE maintains its leading position. As shown in Table 7, BPE achieves an average accuracy of 79.72%, marginally outperforming Simple Average (79.66%) and Weighted Average (79.63%). Consistently, the Friedman test results in Fig. 5 rank BPE first with a score of 3.400, followed by Weighted Average at 3.875.
The statistical analysis presented in Table 8 reveals an interesting pattern distinct from the heterogeneous case. BPE exhibits statistically significant superiority over Stacking ($p\approx 0.0006$) and all DES variants (e.g., DES-AS, KNORA-Union, with $p\ll 0.05$). However, the null hypothesis is not rejected when comparing BPE with Simple Average ($p=0.122$) and Weighted Average ($p=0.072$). This convergence in performance can be attributed to the inherent characteristics of homogeneous pools constructed from unpruned decision trees. Individual unpruned trees are highly sensitive to noise and prone to overfitting, which introduces instability in the local regions used to construct behavioral profiles. While this makes it challenging to significantly outperform robust static fusion methods like Simple Averaging in every instance, BPE’s consistent top-ranking performance across metrics confirms its capability to effectively mitigate these instabilities better than other dynamic selection strategies.
Table 7: Average classification accuracy obtained by BPE and twelve baseline models over 40 datasets in homogeneous ensemble case. The best results are highlighted in bold-face.
Dataset Static DES Proposed
Best Simple Median Rank Stacking Weighted DES-AS KNE KNU LCA MCB RRC BPE
D1 95.73 97.21 97.04 93.44 96.81 97.13 96.93 96.97 97.16 97.24 97.29 97.17 96.97
D2 32.03 33.76 33.50 32.12 32.76 33.85 33.29 33.56 33.59 33.79 33.79 33.74 33.47
D3 62.23 70.71 70.29 68.45 68.42 70.00 68.16 69.19 69.10 69.74 69.61 69.81 69.58
D4 73.78 76.69 76.44 73.29 76.10 76.71 76.67 76.30 76.53 76.50 76.44 76.61 76.77
D5 85.79 89.94 89.71 90.53 89.33 90.01 89.80 89.10 89.68 89.85 89.84 89.83 89.95
D6 68.59 70.20 70.28 60.07 71.93 70.07 70.37 69.96 70.24 70.26 69.85 70.20 70.20
D7 73.74 76.51 76.55 76.29 76.13 76.66 76.40 75.78 76.53 76.48 76.88 76.46 76.88
D8 86.87 89.45 89.19 87.87 89.38 89.55 89.55 88.83 89.53 89.51 89.51 89.47 89.53
D9 77.87 81.94 80.90 85.74 80.32 82.00 82.26 81.61 82.39 82.06 81.94 82.26 82.06
D10 80.36 86.38 83.72 87.54 86.05 86.86 85.65 85.47 85.81 86.44 86.65 86.54 87.75
D11 77.54 81.96 81.98 76.84 82.32 82.42 82.08 81.28 82.21 82.69 81.79 82.69 82.51
D12 70.53 76.03 75.84 76.21 74.94 76.03 74.69 73.96 75.28 75.56 74.68 75.62 76.50
D13 98.64 99.04 99.03 98.21 98.96 99.05 99.03 98.97 99.06 99.06 99.06 99.06 99.06
D14 98.73 99.12 99.12 98.27 99.06 99.13 99.11 99.02 99.13 99.13 99.14 99.14 99.11
D15 87.72 89.42 89.37 80.79 89.44 89.26 89.07 88.95 89.24 89.24 89.19 89.23 89.47
D16 85.44 88.57 88.49 70.30 89.18 88.46 88.20 88.05 88.45 88.39 88.29 88.39 88.63
D17 74.86 81.30 81.19 62.84 81.42 81.30 81.11 79.71 81.29 81.32 81.21 81.34 81.42
D18 79.38 82.60 82.37 63.46 82.70 82.45 81.69 82.05 82.09 82.17 82.20 82.14 82.78
D19 81.65 85.14 84.99 64.04 85.36 85.07 84.87 84.48 84.82 84.89 84.68 84.89 85.23
D20 87.53 89.58 89.40 89.09 89.49 89.59 89.39 89.38 89.47 89.55 89.43 89.55 89.58
D21 87.34 89.79 89.72 79.25 89.80 89.78 89.75 89.32 89.79 89.81 89.76 89.81 89.84
D22 73.40 76.33 75.93 69.47 74.20 76.27 75.47 74.73 75.73 75.93 76.67 76.13 76.67
D23 29.94 37.44 37.31 38.19 31.44 36.38 33.87 37.38 34.75 36.19 36.25 35.81 36.69
D24 99.62 99.98 99.96 99.96 99.97 99.98 99.97 99.97 99.98 99.98 99.98 99.98 99.98
D25 86.96 89.96 89.77 73.16 91.19 90.14 90.04 89.16 90.28 90.25 90.27 90.27 90.37
D26 65.92 68.81 68.95 63.63 70.71 68.49 68.56 68.35 68.65 68.57 68.34 68.56 68.46
D27 83.47 88.47 88.38 84.73 88.44 88.47 88.21 87.05 88.46 88.53 88.41 88.56 88.46
D28 80.33 84.53 84.37 82.12 84.13 84.38 83.98 83.33 84.25 84.40 84.33 84.42 84.30
D29 87.31 93.55 93.55 93.52 93.49 93.52 93.29 91.96 93.47 93.54 93.38 93.53 93.57
D30 10.28 10.73 10.49 10.90 10.09 10.76 10.50 10.73 10.52 10.63 10.63 10.64 10.59
D31 87.53 89.68 89.55 88.36 89.33 89.74 89.77 89.26 89.94 90.00 89.70 89.98 90.33
D32 32.07 38.98 38.45 36.59 38.91 38.98 37.85 38.72 38.48 38.89 38.89 38.89 40.11
D33 47.26 63.76 63.41 62.11 62.26 63.77 62.31 61.16 63.02 63.57 63.57 63.59 63.03
D34 67.54 72.48 71.93 71.35 72.74 72.70 71.59 71.58 72.00 72.40 71.85 72.47 72.48
D35 94.57 98.57 98.11 95.62 98.73 98.69 98.62 97.44 98.68 98.71 98.47 98.68 98.78
D36 89.91 92.01 91.85 91.13 92.05 92.05 91.96 91.76 91.94 91.97 91.76 91.97 92.20
D37 75.81 79.31 78.98 76.84 79.29 79.29 79.20 78.93 79.24 79.25 79.20 79.23 79.39
D38 76.19 87.35 87.01 87.41 86.31 87.18 85.94 84.61 86.30 86.98 87.33 86.81 86.85
D39 96.08 96.88 96.80 88.08 96.86 96.89 96.82 96.81 96.85 96.86 96.88 96.87 96.90
D40 82.05 85.16 84.97 85.15 84.95 85.23 84.99 84.47 85.09 85.11 85.00 85.15 85.20
Average 75.36 79.30 79.03 74.90 78.94 79.28 78.81 78.53 79.02 79.20 79.10 79.20 79.36
Figure 5: Average Friedman ranks of BPE and twelve baseline models in homogeneous ensemble case.
Figure 6: Win-tie-loss distribution of the average classification accuracy of the proposed BPE vs. twelve baseline models over 40 datasets in homogeneous ensemble scenario. Each full line illustrates the critical value $n_{c}=\{24.05,25.20,27.36\}$ considering confidence level of $\alpha=\{0.10,0.05,0.01\}$.
Table 8: Wilcoxon signed-rank test results for comparing BPE with twelve baseline models in homogeneous ensemble case.
Ensemble type Comparison $^{+} RR^{-}$ Hypothesis $p$-value
Static BPE vs. Best Single 820.0 0.0 Rejected for BPE at 5% 0.0000**
Static BPE vs. Median Avg 664.0 156.0 Rejected for BPE at 5% 0.0004**
Static BPE vs. Rank Avg 752.0 68.0 Rejected for BPE at 5% 0.0000**
Static BPE vs. Simple Avg 477.0 264.0 Not Rejected 0.1225
Static BPE vs. Stacking 658.0 162.0 Rejected for BPE at 5% 0.0006**
Static BPE vs. Weighted Avg 546.0 274.0 Not Rejected 0.0682
DES BPE vs. DES-AS 776.0 44.0 Rejected for BPE at 5% 0.0000**
DES BPE vs. KNE 753.0 27.0 Rejected for BPE at 5% 0.0000**
DES BPE vs. KNU 689.0 91.0 Rejected for BPE at 5% 0.0000**
DES BPE vs. LCA 539.0 202.0 Rejected for BPE at 5% 0.0145**
DES BPE vs. MCB 586.0 117.0 Rejected for BPE at 5% 0.0004**
DES BPE vs. RRC 536.0 205.0 Rejected for BPE at 5% 0.0164**
** Close to the $p$-value means that statistical differences are found with $\alpha=0.05$.
6 Discussion and Future Work
The BPE framework fundamentally represents a paradigm shift. It transitions from the traditional ensemble approach, which bases integration on the performance disparity among different models, to a novel perspective that bases integration on the variation of a single model’s performance relative to its own internal behavioral profile.
In this work, the proposed BPE-Entropy algorithm should be viewed as a minimal and representative instantiation of this general framework, rather than an attempt to exhaustively optimize the metric. The use of negative entropy and Gaussian perturbations is deliberately chosen for their conceptual simplicity and theoretical interpretability, allowing the core principles of the BPE framework to be clearly isolated and evaluated.
Admittedly, consistent with the “No Free Lunch” theorem in machine learning, the BPE algorithm is not universally optimal across all scenarios and encounters specific challenges and difficult conditions. In particular, as a nascent ensemble perspective, it remains in an early stage of development. Based on the intrinsic characteristics of this framework, we propose the following potential directions for future research:
  • 1.
    Metrics for Behavioral Profiling: A critical question remains: what should serve as the model’s behavioral profile? In this study, the BPE framework utilizes the entropy of the predicted output as the profiling metric. However, whether entropy can fully and accurately characterize a model’s behavioral properties remains an open question. We also explored using the margin between the Top-1 and Top-2 prediction probabilities as an alternative profile, which outperformed entropy in certain scenarios. Consequently, identifying a metric that accurately and comprehensively describes a model’s behavioral characteristics constitutes a significant problem for future investigation.
  • 2.
    Construction of Behavioral Profiles: Currently, we construct the behavioral profile through a simple noise injection strategy on the training set. This is a relatively rudimentary implementation, as the primary focus of this work is to highlight the shift in ensemble perspective rather than to pursue the ultimate performance limit. Therefore, developing more sophisticated methods for constructing behavioral profiles will be key to raising the performance ceiling. A distinct advantage of the BPE framework over methods like DES and Stacking is that it does not require a validation set. This characteristic allows for various data augmentation techniques on the training set without the need for ground truth labels, thereby mitigating the limitations imposed by small-sample environments compared to traditional approaches.
  • 3.
    Integration of Perspectives: Traditional ensemble methods focus on evaluating how well a model performs relative to other models, whereas the BPE framework focuses on evaluating how well a model performs relative to itself. These two ideologies are not mutually exclusive but rather complementary. Thus, investigating how to effectively integrate these two perspectives is a highly valuable research direction. Preliminary experiments conducted in this study suggest that fusing the weights derived from both ideologies can indeed achieve a higher performance upper bound in many scenarios.